# The cognitive triple-slit experiment

### v2

@@ 1 Introduction
Figure 1: A schematic representation of the two-slit experiment in the case of classical and quantum entities. In the classical case (top right), the pattern on the screen can be deduced from a uniform average of the two patterns when the two slits are opened one at a time. In the quantum case (bottom right), the fringe pattern cannot be deduced from those obtained when only one slit remains open at a time.
@@
Cognitive equivalents of the double-slit experiment also exist [11]. Let us consider an important example. In the 1980s, psychologist James Hampton conducted an experiment in which 40 students were presented with $24$ exemplars of Food and asked to what extent they thought they were related (in terms of membership and typicality) to the concepts: (a) Fruit, (b) Vegetable, and (c) Fruit or vegetable [12]. The $24$ exemplars of Food played the same role as the different locations where the trace of an impact can be observed on the detection screen of a double-slit experiment, with the two abstract concepts Fruit and Vegetable playing the role of the two slits. More precisely, choosing a typical exemplar of Fruit is here equivalent to choosing a good example of an impact of an entity passing through the first slit. Similarly, selecting a typical exemplar of Vegetable is equivalent to providing a good example of the impact of an entity passing through the second slit. Finally, selecting a typical exemplar of Fruit or vegetable is equivalent to providing a good example of an impact of an entity passing through both slits, i.e., through the first slit or the second slit.
@@
An expectation based on classical prejudice would tell us that when faced with question (c), about the combination Fruit or vegetable, the probability of selecting a given exemplar of Food would be the uniform average of the probabilities describing the situations of questions (a) and (b), but this is not what Hampton’s data revealed. Similar to the constructive and destructive interference effects in the impact statistics of typical double-slit experiments, overextensions and underextensions of the probabilities were observed, and when Hampton’s data were represented in a quantum-like way, using two Gaussian wave functions and their superpositions, an interesting interference-like figure was revealed, reminiscent of those obtained in the phenomena of birefringence [11]; see Figure 2 see also Figure 10, for a one-dimensional modelization of Hampton’s data).
@@
Figure 2: (color online) The first two figures on the left represent, respectively, the experimental probabilities of selecting a typical exemplar of Fruit, with Apple being the most frequently chosen, and the experimental probabilities of selecting a typical exemplar of Vegetable, with Broccoli being the most frequently chosen. The figure at the top right represents the probabilities obtained by considering a uniform average of the probabilities for Fruit and for Vegetable. This average, typical for classical entities, differs greatly from the data observed by Hampton, when the conceptual combination Fruit or vegetable was considered. The figure at the bottom right represents the experimental probabilities of selecting a typical exemplar of Fruit or vegetable, modeled using the quantum superposition principle. It exhibits interference phenomena similar to those of birefringence. For more information on how these curves were constructed, see Section 6 and [11].
@@
A possible interpretation of the overextension and underextension effects obtained in Hampton’s experiment is that they would correspond to regions minimizing (respectively, maximizing) the doubt of the participants as to how to classify a given exemplar. Let us explain this in some detail. When the students were asked about a typical exemplar of Fruit, the most frequent answer (the one receiving the higher rating) was Apple, and when asked about a typical exemplar of Vegetable, the most frequent answer was Broccoli. On the other hand, Mushroom, although considered to be more representative of Vegetable than of Fruit, it wasn’t considered a typical exemplar of either, in the sense that the associated probabilities were very small compared to those associated with Apple and Broccoli. But when the students were asked about a typical exemplar of the conceptual combination Fruit or vegetable, it was Mushroom that received the higher score, with a probability that was almost double compared to the uniform average of the probabilities related to only Fruit and only Vegetable (overextension effect). This deviation can be explained by observing that the disjunction Fruit or vegetable is a new emerging concept, not reducible, in regard to its meaning, to the meanings of its components taken individually. More precisely, Fruit or vegetable, in addition to its pure logical meaning, also conveys the meaning of a doubt as to whether a given exemplar of Food can be classified as Fruit or as Vegetable, and since Mushroom does not classify well neither as Fruit nor as Vegetable, it will be an excellent exemplar to be classified as Fruit or vegetable, hence the overextension effect. Similarly, an exemplar like Elderberry, since it is typical of Fruit and not typical of Vegetable, and is scored accordingly when answering questions (a) and (b), it is not a good example of a situation of doubt as to whether it belongs to one of these two categories. So, it will receive a score lower than the uniform average of the probabilities related to only Fruit and only Vegetable, hence the underextension effect.
@@
If we follow this line of reasoning, we can try to provide a cognitive-like interpretation of the double-slit experiment by claiming that the fringes of higher (lower) intensity in the interference pattern correspond to regions that maximize (minimize) doubt about the slits from which the impact came from. In other words, an impact-rich fringe would be a region on the screen that makes it difficult to guess which slit the quantum entity passed through, while it would be exactly the opposite for the regions on the screen with a low density of impacts [11]. Such an interpretation is part of the so-called conceptuality interpretation of quantum mechanics, first proposed by Diederik Aerts in 2009 [11] and further developed in his Brussels group [13, 14, 15, 16, 17, 18, 19, 20]. According to it, the behavior of quantum entities appears to be bizarre only if observed on the basis of the prejudice that they are spatiotemporal in nature, for example corpuscles, waves, or even fields. However, if we hypothesize that their nature is conceptual, and that quantum measurements are meaning-driven processes where the measuring apparatuses select outcomes similarly to how cognitive entities respond to questions, then many of the quantum conundrums disappear. Of course, human concepts and the conceptual entities of the microworld should not be confused, in the same way that, for example, electromagnetic and acoustic waves should not be confused, although they share the same undulatory nature.
@@
In this regard, it is important to know that, in the last two decades, a field of research called quantum cognition has emerged, using quantum mechanics, and especially its mathematics, to find new tools for explaining and modeling human cognition, especially how meaning is assigned to concepts and their combinations, and how decision-making processes occur depending on context. It is because of the success of this quantum approach to human modeling that the speculative hypothesis at the basis of the conceptuality interpretation subsequently emerged, according to which this success was not accidental, but due to the fact that quantumness and conceptuality are notions that point to a same nature. Indeed, by using the quantum mathematical formalism and the notions that this formalism brings to bear (in particular, potentiality and context-dependence), quantum cognition was able to bring to the fore, in the psychological laboratories, all those effects that are typical of quantum entities, such as superposition and interference effects, entanglement, complementarity and indistinguishability [19, 20, 21, 22, 23, 24, 25, 26, 27].
@@ 2 The quantum triple-slit experiment
For corpuscular entities, the average of the two single-slit situations reproduce the double-slit situation. More precisely, denoting $P_{\rm cl}(x|12)$ the (classical) probability of observing an impact at point $x$ of the detection screen, when both slits are open, we can always write it as the uniform average
$P_{\rm cl}(x|12)=\tfrac{1}{2}P_{\rm cl}(x|1)+\tfrac{1}{2}P_{\rm cl}(x|2)$ (1)
where we have assumed that the probabilities of passing through slit-$1$ and slit-$2$ are the same, and $P_{\rm cl}(x|1)$ and $P_{\rm cl}(x|2)$ are the probabilities of observing an impact at point $x$ when only slit-$1$ and slit-$2$ are open, respectively. As mentioned in the previous section, the above average does not hold anymore if the entities are quantum. More precisely, if $\psi(x|1)$ and $\psi(x|2)$ are the wave functions at the detection screen (assumed here to be orthogonal unit vectors), describing the situations where the quantum entity passed through slit-$1$ and slit-$2$ only, respectively, then we know, according to the Born rule, that the associated quantum probabilities are given by the squared modules $P(x|1)=|\psi(x|1)|^{2}$ and $P(x|2)=|\psi(x|2)|^{2}$. When both slits are open, we can apply the prescription saying that when there are alternatives (slit-$1$ or slit-$2$), the probability amplitude, $\psi(x|12)$, is obtained from the normalized sum of the probability amplitudes associated with the alternatives considered separately:
$\psi(x|12)=\tfrac{1}{\sqrt{2}}\psi(x|1)+\tfrac{1}{\sqrt{2}}\psi(x|2)$ (2)
Therefore, the associated probability is:
$P(x|12)=|\psi(x|12)|^{2}=\overline{P}(x|12)+I(x|12)$ (3)
where we have defined the uniform (classical) average $\overline{P}(x|12)\equiv\tfrac{1}{2}P(x|1)+\tfrac{1}{2}P(x|2)$, and the two-slit interference contribution $I(x|12)\equiv\Re\,[\psi(x|1)^{*}\psi(x|2)]$, which accounts for the overextension and underextension effects ($\Re$ denotes the real part of a complex number).
With obvious notation, in the situation with three slits one has to consider the superposition:
@@
$P(x|123)=|\psi(x|123)|^{2}=\overline{P}(x|123)+\tfrac{2}{3}\left[I(x|12)+I(x|23)+I(x|13)\right]$ (5)
where we have defined the (classical) average $\overline{P}(x|123)\equiv\tfrac{1}{3}P(x|1)+\tfrac{1}{3}P(x|2)+\tfrac{1}{3}P(x|3)$, and the two-slit interference contributions $I(x|23)\equiv\Re\,[\psi(x|2)^{*}\psi(x|3)]$, $I(x|13)\equiv\Re\,[\psi(x|1)^{*}\psi(x|3)]$. Inserting (3) in (5), and same for the other two interference terms, after due simplification one gets:
@@
called the Sorkin parameter [28], one obtains the condition: $\epsilon(x|123)=0$, for all $x$. So, the Sorkin parameter, an expression consisting only of experimentally accessible probabilities, quantifies possible deviations of the three-slit interference contribution, $I(x|123)\equiv P(x|123)-\overline{P}(x|123)$, from contributions only coming from second-order two-slit interference effects, as is clear that it can also be written
@@ 3 Third-order interference effects
As first pointed out by Rafael Sorkin [28], the Born rule being quadratic in the wave function it does not allow for genuine third-order interference contributions [29, 30]. From this follows that the Sorkin parameter (7) is zero, which is something that can be experimentally tested. The first experiments to measure the Sorkin parameter were performed in 2009 by Urbasi Sinha and colleagues [31, 32, 33]. Their result was that, within their error margins, the nullity of the Sorkin parameter had been verified, and in this sense the prediction of the Born rule confirmed. More precisely, the magnitude of a third-order interference contribution was shown to be less than $10^{-2}$ of the magnitude of the expected second-order interference contributions. Note that any setting that allows for the presence of three mutually exclusive paths can be used to test the nullity of the Sorkin parameter. This was the situation of the experiment performed by Söllner and colleagues [34] with a three-path photonic interferometer, which allowed them to improve by about an order of magnitude the best previous experiments.
@@
Up to this point, it was believed that Sorkin’s initial analysis and subsequent experiments were to be interpreted as confirming the peculiarity of the Born rule in not predicting genuine third-order interference contributions. But things were not as they appeared. As pointed out in [35], it was incorrect to assume that it is always possible to write the wave function on the detection screen, when all three slits are open, as the superposition (4) of contributions due only to the three alternatives in play, i.e., to passing through only one of the three slits. One way to understand this is to observe that, due to the phenomenon of the spreading of the wave-packet, it is not possible to decompose a wave-function on the screen as a superposition of wave-functions uniquely associated with the single-slit situations. In fact, during its evolution in time, the support of any wave function will inevitably always extend to touch with its tails all the slits. This can be expressed in a more specific way by considering a Feynman path-integral approach [36], which involves integration over all possible paths, including those passing through several slits (forming loops for example) before reaching the detection screen [37], and this is another way of saying that the wave function is not uniquely decomposable as a superposition of contributions attributable to individual slits. While these looped paths contribute less to the overall probability than paths closer to classical trajectories, they cannot be neglected. In other words, contrary to the initial belief, a non-zero value of the Sorkin parameter was not to be considered as a violation of the Born rule [35, 38, 39, 40]. To explain it yet another way, Yabuki emphasized that when we close slits we modify the Hamiltonian. Thus, in a double-slit problem the solutions with only one slit open are not directly related to the solutions with both slits open, since they correspond to different physical contexts, so the superposition principle cannot be applied naively [37]. And of course the same is true for a three-slit experiment.
@@
Figure 3: Three examples of exotic paths that cross all three slits. See the analysis in [37].
It should be emphasized that even in the two-slit case there are exotic paths that pass through both slits and are contained in the second-order interference contributions. This means that a non-zero Sorkin parameter accounts for complex paths that interact with all three slits. If they usually constitute a minimal correction [39], they become more relevant as the operating wavelength increases with respect to the slit spacing. This is because the longer the wavelength, the greater the overlap between the wave functions of the individual slits, which gives more weight to the non-classical paths interacting with all the slits. This explains why the 2018 experiment by Govindaraj Rengaraj et al., that measured a non-zero Sorkin parameter, worked in the microwave regime [43]. So, in quantum mechanics, when we are in frequency regimes (or experimental designs) where the exotic paths contributions are not relevant, there are no genuine third-order interference effects, but when we move from the physical to the psychological laboratory, things may change, and it is one of the aims of this article to investigate this possibility.
@@
We are not aware of any tests for possible violations of the nullity of the Sorkin parameter in experiments with combinations of three concepts, and it is certainly our intention, in the near future, to carry out such an experiment. However, we can certainly reason as follows. The three two-word disjunctions, Fruit or vegetable, Fruit or flavoring and Vegetable or flavoring, express different situations of doubt about the classification of a given exemplar of Food. Imagine that an exemplar cannot be conveniently classified as either Fruit, Vegetable, or Flavoring. This will lead the corresponding two-word conjunctions receiving higher scores, i.e., being overextended, precisely because they express a situation of doubt. However, the three-word conjunction, Fruit or vegetable or flavoring, also conveys a meaning of doubt about how to classify the exemplar in question, and undoubtedly does so differently from the two-word conjunctions, since it explicitly contains all three words. The question that naturally arises is whether this three-word (three-slit-like) expression of doubt allows, in real experiments, meanings to be derived that are in no way deducible from two-word (two-slit-like) expressions of doubt. If this is the case, the equality $\epsilon(x|123)=0$ is going to be violated. But will it be a weak violation, as is usually the case in quantum mechanics, or a strong one? The cognitive test we have conducted, which is described in the next section and simulates single-slit, double-slit and triple-slit experiments, asking participants to place impact points directly on the equivalent of a detection screen, will allow us to offer an initial response to this question.

### v3

@@ 1 Introduction
Figure 1: A schematic representation of the two-slit experiment in the case of classical (top right) and quantum (bottom right) entities.
@@
Cognitive analogues of the double-slit experiment have also been proposed [11]. We will examine a significant example in detail, as this will clarify the design of the cognitive experiment presented in Section 4. In the 1980s, psychologist James Hampton conducted an experiment in which 40 students were presented with $24$ exemplars of Food and asked to what extent they thought they were related (in terms of membership and typicality) to the concepts: (a) Fruit, (b) Vegetable, and (c) Fruit or vegetable [12]. The $24$ exemplars of Food played the same role as the different locations where the trace of an impact can be observed on the detection screen of a double-slit experiment, with the two abstract concepts Fruit and Vegetable playing the role of the two slits. More precisely, choosing a typical exemplar of Fruit is here equivalent to choosing a good example of an impact of an entity passing through the first slit. Similarly, selecting a typical exemplar of Vegetable is equivalent to providing a good example of the impact of an entity passing through the second slit. Finally, selecting a typical exemplar of Fruit or vegetable is equivalent to providing a good example of an impact of an entity passing through both slits, i.e., through the first slit or the second slit.
@@
An expectation based on classical prejudice would tell us that when faced with question (c), about the combination Fruit or vegetable, the probability of selecting a given exemplar of Food would be the uniform average of the probabilities describing the situations of questions (a) and (b), but this is not what Hampton’s data revealed. Similar to the constructive and destructive interference effects in the impact statistics of typical double-slit experiments, overextensions and underextensions of the probabilities were observed, and when Hampton’s data were represented in a quantum-like way, a complex interference-like pattern was revealed, reminiscent of those obtained in the phenomena of birefringence [11]; see Figure 2.
@@
Figure 2: (color online) On the left are the experimental probabilities of selecting a typical exemplar of Fruit and of Vegetable, with Apple and Broccoli being the most frequently chosen, respectively. The top right figure shows their uniform average, which differs greatly from the experimental probabilities of selecting a typical exemplar of Fruit or vegetable, modeled using the quantum superposition principle (bottom right figure) and exhibiting birefringence-like interference phenomena [11].
More precisely, to obtain this pattern, Aerts represented the concepts Fruits, Vegetables and Fruits or vegetables by unit vectors $|F\rangle$, $|V\rangle$ and $|FV\rangle={1\over\sqrt{2}}(|F\rangle+|V\rangle)$, respectively. He then chose them so that the wave functions $\langle x|F\rangle$ and $\langle x|V\rangle$ would be two-dimensional, with a Gaussian real part. Placing all the exemplars on a two-dimensional plane, in such a way that, for their coordinates, the quantum probabilities $P_{F}(x)=|\langle x|F\rangle|^{2}$, $P_{V}(x)=|\langle x|V\rangle|^{2}$ and $P_{FV}(x)=|\langle x|FV\rangle|^{2}$ properly modeled the experimental probabilities, he then obtained the pattern shown in Figure 2 [11].
Note that the use of Gaussians requires at least two dimensions to distribute the different datapoints. This does not mean that it is not possible to also describe Hampton’s data using a one-dimensional screen of exemplars. For this, one has to position them in such a way that two clear peaks appear for the two single-slit experiments relating to the concepts Fruits and Vegetables, centered on the exemplars that obtained the highest score. Obviously, this requires giving up perfectly symmetrical Gaussian distributions. Once this is done, one can deduce the two-slit curve for the combination Fruits or Vegetables, as shown in Figure 3.
Figure 3: (color online) A continuous approximation of the discrete single-slit functions $P_{F}(x)$ (purple color) and $P_{V}(x)$ (green color), and corresponding double-slit function $P_{FV}(x)$ (brown color), which is the one-dimensional equivalent of the two-dimensional pattern exhibited in Figure 2.
@@
A possible interpretation of the quantum-like [^1] As Figure 3 shows, the inequality $P_{FV}(x)\geq\max\{P_{F}(x),P_{V}(x)\}$ is violated by many exemplars. Therefore, Hampton’s disjunction effects cannot be consistently modeled using a single classical probability space. The same will be true for the probabilities obtained in our experiment. See for example the analysis in [13]. overextension and underextension effects obtained in Hampton’s experiment is that they would correspond to regions minimizing (respectively, maximizing) the doubt of the participants as to how to classify a given exemplar. Let us explain this in some detail. When the students were asked about a typical exemplar of Fruit, the most frequent answer (the one receiving the higher rating) was Apple, and when asked about a typical exemplar of Vegetable, it was Broccoli. On the other hand, Mushroom, although considered to be more representative of Vegetable than of Fruit, it wasn’t considered a typical exemplar of either, in the sense that the associated probabilities were very small compared to those associated with Apple and Broccoli. But when the students were asked about a typical exemplar of the conceptual combination Fruit or vegetable, it was Mushroom that received the higher score, with a probability that was almost double compared to the uniform average of the probabilities related to only Fruit and only Vegetable (overextension effect). This deviation can be explained by observing that the disjunction Fruit or vegetable is a new emerging concept, not reducible, in regard to its meaning, to the meanings of its components taken individually. More precisely, Fruit or vegetable, in addition to its pure logical meaning, also conveys the meaning of a doubt as to whether a given exemplar of Food can be classified as Fruit or as Vegetable, and since Mushroom does not classify well neither as Fruit nor as Vegetable, it will be an excellent exemplar to be classified as Fruit or vegetable, hence the overextension effect. Similarly, an exemplar like Elderberry, since it is typical of Fruit and not typical of Vegetable, and is scored accordingly when answering questions (a) and (b), it is not a good example of a situation of doubt as to whether it belongs to one of these two categories. So, it will receive a score lower than the uniform average of the probabilities related to only Fruit and only Vegetable, hence the underextension effect.
@@
If we follow this line of reasoning, we can try to provide a cognitive-like interpretation of the double-slit experiment by claiming that the fringes of higher (lower) intensity in the interference pattern correspond to regions that maximize (minimize) doubt about the slits from which the impact came from. In other words, an impact-rich fringe would be a region on the screen that makes it difficult to guess which slit the quantum entity passed through, while it would be exactly the opposite for the regions on the screen with a low density of impacts [11].
This analogy with Hampton’s experiment requires interpreting Fruit and Vegetable as two possible states of the concept Food. The conceptual combination The entity passes through slit-1 is then the analogue of the conceptual combination The food is a fruit, the conceptual combination The entity passes through slit-2 is the analogue of The food is a vegetable, and the conceptual combination The entity passes through slit-1 or slit-2 is the analogue of The food is a fruit or vegetable. The selected exemplars also correspond to possible states of the Food conceptual entity, such as the state The food is an apple. These are the analogues of the states corresponding to the different locations where the quantum entity can be detected on the screen, such as The entity is located at position $x=0$. Hence, instead of saying “the entity passing through slit-1 (respectively, slit-2) is detected in position xx,” we should say “position $x$ is a good example of the entity passing through slit-1 (respectively, slit-2).” Similarly, instead of saying “the entity passing through slit-1 or slit-2 is detected in position xx,” we should say “position $x$ is a good example of the entity passing through slit-1 or slit-2.”
Following the above analogy, which is rooted in the conceptuality interpretation of quantum mechanics (first proposed by Diederik Aerts in 2009 [11] and further developed by his Brussels group [14, 15, 16, 17, 18, 19, 20, 21]), the detection screen must be viewed as a cognitive structure sensitive to the meaning carried by the entities interacting with it. Outcomes are thus selected in a manner analogous to how cognitive entities respond to questions. In other words, within this interpretation, quantum entities are assumed to be conceptual rather than spatiotemporal, and this shift in perspective explains their apparently paradoxical behavior (human concepts and the conceptual entities of the microworld should however not be conflated, just as electromagnetic and acoustic waves should not be confused, despite both sharing an undulatory nature).
In the double-slit experiment, when the state expresses a lack of knowledge, i.e., incomplete information about the slit through which the entity passes, this incompleteness is associated, in the conceptual domain where meaning-driven interactions occur, with a genuine new element of reality. Hence, in the quantum formalism, the situation is described by a superposition state rather than a mixture. The central fringe then becomes the region where the best answers to the question posed are located, in the sense of providing the most faithful exemplifications of an ontic, not merely epistemic, condition of indeterminacy.
@@
It is worth mentioning that in the last two decades a field of research called quantum cognition has emerged, using quantum mechanics, and especially its mathematics, to find new tools for explaining and modeling human cognition, especially how meaning is assigned to concepts and their combinations, and how decision-making processes occur depending on context. It is because of the success of this quantum approach to human modeling that the speculative hypothesis at the basis of the conceptuality interpretation subsequently emerged, according to which this success was not accidental, but due to the fact that quantumness and conceptuality are notions that point to a same nature. Indeed, by using the quantum formalism and the notions that it brings to bear (in particular, potentiality and context-dependence), quantum cognition was able to bring to the fore, in the psychological laboratories, all those effects that are typical of quantum entities, such as superposition and interference effects, entanglement, complementarity and indistinguishability [20, 21, 22, 23, 24, 25, 26, 27, 28].
@@ 2 The quantum triple-slit experiment
For classical, corpuscular entities, the average of the two single-slit situations reproduce the double-slit situation. More precisely, denoting $P_{\rm cl}(x|12)$ the (classical) probability of observing an impact at point $x$ of the detection screen, when both slits are open, we can always write it as the uniform average $P_{\rm cl}(x|12)=\tfrac{1}{2}P_{\rm cl}(x|1)+\tfrac{1}{2}P_{\rm cl}(x|2)$, where we have assumed that the probabilities of passing through slit-$1$ and slit-$2$ are the same, and $P_{\rm cl}(x|1)$ and $P_{\rm cl}(x|2)$ are the probabilities of observing an impact at point $x$ when only slit-$1$ and slit-$2$ are open, respectively. As mentioned in the previous section, the above average does not hold anymore if the entities are quantum. More precisely, if $\psi(x|1)$ and $\psi(x|2)$ are the wave functions at the detection screen, assumed here to be orthogonal unit vectors, [^2] The orthogonality of the one-slit wave functions can be reasonably assumed because the slits correspond to spatially disjoint regions. Therefore, the wave functions can be considered orthogonal at the slit plate level, since their supports do not overlap. Since unitary evolution preserves orthogonality, the wave functions will also remain orthogonal at the detection screen level, even though they overlap at the screen and produce interference contributions. (Orthogonality in Hilbert space does not forbid overlap in configuration space.) describing the situations where the quantum entity passed through slit-$1$ and slit-$2$ only, respectively, then we know, according to the Born rule, that the associated quantum probabilities are given by the squared modules $P(x|1)=|\psi(x|1)|^{2}$ and $P(x|2)=|\psi(x|2)|^{2}$. When both slits are open, we can apply the prescription saying that when there are alternatives (slit-$1$ or slit-$2$), the probability amplitude, $\psi(x|12)$, is obtained from the normalized sum of the probability amplitudes associated with the alternatives considered separately: $\psi(x|12)=\tfrac{1}{\sqrt{2}}\psi(x|1)+\tfrac{1}{\sqrt{2}}\psi(x|2)$. Therefore, the associated probability is:
$P(x|12)$ $=$ $|\psi(x|12)|^{2}=|\tfrac{1}{\sqrt{2}}\psi(x|1)+\tfrac{1}{\sqrt{2}}\psi(x|2)|^{2}$ (1)
$=$ $\tfrac{1}{2}P(x|1)+\tfrac{1}{2}P(x|2)+\Re\,[\psi(x|1)^{*}\psi(x|2)]=\overline{P}(x|12)+I(x|12)$
where we have defined the uniform average $\overline{P}(x|12)\equiv\tfrac{1}{2}P(x|1)+\tfrac{1}{2}P(x|2)$, and the two-slit interference contribution $I(x|12)\equiv\Re\,[\psi(x|1)^{*}\psi(x|2)]$, which accounts for the overextension and underextension effects.
In the situation with three slits one has to consider the superposition:
@@
$P(x|123)$ $=$ $|\psi(x|123)|^{2}=\tfrac{1}{3}P(x|1)+\tfrac{1}{3}P(x|2)+\tfrac{1}{3}P(x|3)$ (3)
$+$ $\tfrac{2}{3}\Re\,[\psi(x|1)^{*}\psi(x|2)]+\tfrac{2}{3}\Re\,[\psi(x|2)^{*}\psi(x|3)]+\tfrac{2}{3}\Re\,[\psi(x|1)^{*}\psi(x|3)]$
$=$ $\overline{P}(x|123)+\tfrac{2}{3}\left[I(x|12)+I(x|23)+I(x|13)\right]$
where we have defined the uniform average $\overline{P}(x|123)\equiv\tfrac{1}{3}\sum_{i=1}^{3}P(x|i)$, and the two-slit interference contributions $I(x|12)\equiv\Re\,[\psi(x|1)^{*}\psi(x|2)]$, $I(x|23)\equiv\Re\,[\psi(x|2)^{*}\psi(x|3)]$ and $I(x|13)\equiv\Re\,[\psi(x|1)^{*}\psi(x|3)]$. By definition of the latter, we can also write $I(x|12)=P(x|12)-\overline{P}(x|12)$, $I(x|23)=P(x|23)-\overline{P}(x|23)$, and $I(x|13)=P(x|13)-\overline{P}(x|13)$. Inserting these three expressions in (3), then observing that, by definition of uniform averages, we also have
$\overline{P}(x|123)-\tfrac{2}{3}\left[\overline{P}(x|12)+\overline{P}(x|23)+\overline{P}(x|13)\right]=-\tfrac{1}{3}\overline{P}(x|123)$ (4)
one obtains the identity:
@@
called the Sorkin parameter [29], one obtains the condition: $\epsilon(x|123)=0$, for all $x$. So, the Sorkin parameter, an expression consisting only of experimentally accessible probabilities, quantifies possible deviations of the three-slit interference contribution from contributions only coming from second-order two-slit interference effects. Indeed, inserting $I(x|123)\equiv P(x|123)-\overline{P}(x|123)$ in (3), it can also be written
@@ 3 Third-order interference effects
As first pointed out by Rafael Sorkin [29], the Born rule being quadratic in the wave function it does not allow for genuine third-order interference contributions [30, 31]. From this follows that the Sorkin parameter (6) is zero, which is something that can be experimentally tested. The first experiments to measure the Sorkin parameter were performed in 2009 by Urbasi Sinha and colleagues [32, 33, 34]. Their result was that, within their error margins, the nullity of the Sorkin parameter had been verified, and in this sense the prediction of the Born rule had been confirmed. More precisely, the magnitude of a third-order interference contribution was shown to be less than $10^{-2}$ of the magnitude of the expected second-order interference contributions.
Note that any setting that allows for the presence of three mutually exclusive paths can be used to test the nullity of the Sorkin parameter. This was the situation of the experiment performed by Söllner and colleagues [35] with a three-path photonic interferometer, which allowed them to improve by about an order of magnitude the best previous experiments.
@@
Up to this point, it was believed that Sorkin’s initial analysis and subsequent experiments were to be interpreted as confirming the peculiarity of the Born rule in not predicting genuine third-order interference contributions. But things were not as they appeared. As pointed out in [36], it was incorrect to assume that it is always possible to write the wave function on the detection screen, when all three slits are open, as the superposition (2) of contributions due only to the three alternatives in play, i.e., to passing through only one of the three slits. One way to understand this is to observe that, due to the phenomenon of the spreading of the wave-packet, it is not possible to decompose a wave-function on the screen as a superposition of wave-functions uniquely associated with the single-slit situations. In fact, during its evolution in time, the support of any wave function will inevitably always extend to touch with its tails all the slits. This can be expressed in a more specific way by considering a Feynman path-integral approach [37], which involves integration over all possible paths, including those passing through several slits (forming loops for example; see Fig. 4) before reaching the detection screen [38], and this is another way of saying that the wave function is not uniquely decomposable as a superposition of contributions attributable to individual slits. While these looped paths contribute less to the overall probability than paths closer to classical trajectories, they cannot be neglected. In other words, contrary to the initial belief, a non-zero value of the Sorkin parameter was not to be considered as a violation of the Born rule [36, 39, 40, 41]. To explain it yet another way, Yabuki emphasized that when we close slits we modify the Hamiltonian. Thus, in a double-slit problem the solutions with only one slit open are not directly related to the solutions with both slits open, since they correspond to different physical contexts, so the superposition principle cannot be applied naively [38]. And of course the same is true for a three-slit experiment.
Figure 4: Examples of exotic paths that cross all three slits. See the analysis in [38].
@@
It should be emphasized that even in the two-slit case there are exotic paths that pass through both slits and are contained in the second-order interference contributions. This means that a non-zero Sorkin parameter only accounts for complex paths that interact with all three slits. If they usually constitute a minimal correction [40], they become more relevant as the operating wavelength increases with respect to the slit spacing. This is because the longer the wavelength, the greater the overlap between the wave functions of the individual slits, which gives more weight to the non-classical paths interacting with all the slits. This explains why the 2018 experiment by Govindaraj Rengaraj et al., that measured a non-zero Sorkin parameter, worked in the microwave regime [44]. So, in quantum mechanics, when we are in frequency regimes (or experimental designs) where the exotic paths contributions are not relevant, there are no genuine third-order interference effects, but when we move from the physical to the psychological laboratory, things may change, and it is one of the aims of this article to investigate this possibility.
@@
We are not aware of any tests for possible violations of the nullity of the Sorkin parameter in experiments with combinations of three concepts, and it is certainly our intention, in the near future, to carry out such an experiment. However, we can certainly reason as follows. The three two-word disjunctions, Fruit or vegetable, Fruit or flavoring and Vegetable or flavoring, express different situations of doubt about the classification of a given exemplar of Food. Imagine that an exemplar cannot be conveniently classified as either Fruit, Vegetable, or Flavoring. This will lead the corresponding two-word conjunctions receiving higher scores, i.e., being overextended, precisely because they express a situation of doubt. However, the three-word conjunction, Fruit or vegetable or flavoring, also conveys a meaning of doubt about how to classify the exemplar in question, and undoubtedly does so differently from the two-word conjunctions, since it explicitly contains all three words. The question that naturally arises is whether this three-word (three-slit-like) expression of doubt allows, in real experiments, meanings to emerge that are in no way deducible from two-word (two-slit-like) expressions of doubt. If this is the case, the equality $\epsilon(x|123)=0$ is going to be violated. But will it be a weak violation, as is usually the case in quantum mechanics, or a strong one? The cognitive test we have conducted, which is described in the next section and simulates single-slit, double-slit and triple-slit experiments, asking participants to place impact points directly on the equivalent of a detection screen, will allow us to offer an initial response to this question.